solue the following equation x-2 X-6 X+3 x+5
Answer #1


Son2 na +3n-10 n + 3nt 18 20 X-2 X-6 x+5 x+3 (M-2)(n+5) = (2-6) (243) 2 n- tn- 2n-10 = n +32-64-18 n-3n-18 2 on +3n-10 2 2 on